OpenStudy (latoya):

find the number. four times a number, added to 7 is 9

OpenStudy (anonymous):

(9-7)/4 = x

OpenStudy (latoya):

now i have to find what x is

OpenStudy (anonymous):

Fo' shizzle.

OpenStudy (aroub):

Let the number be x 4x+7=9 4x=2 x=1/2

OpenStudy (latoya):

the sum of-5, -3 and a number is -12. find the number for x.